determine the particular solution of the differential equations y''+9y=e^xcosx, using the method of undetermined coefficient


1
Expert's answer
2021-01-07T16:03:49-0500

"y''+9y=e^xcosx"


we will find particular solution in the form:

"y_p=Ae^xcosx+Be^xsinx"


"y_p'=Ae^xcosx-Ae^xsinx+Be^xsinx+Be^xcosx="

"=(A+B)e^xcosx+(B-A)e^xsinx"


"y_p''=(A+B)e^xcosx-(A+B)e^xsinx+"

"+(B-A)e^xsinx+(B-A)e^xcosx="

"=2Be^xcosx-2Ae^xsinx"


After substitution into equation:

"2Be^xcosx-2Ae^xsinx+9Ae^xcosx+9Be^xsinx="

"=(2B+9A)e^xcosx+(9B-2A)e^xsinx=e^xcosx"

We have a system of equations:

"2B+9A=1"

"9B-2A=0"


"A=\frac{9B}{2}"

"2B+9*\frac{9B}{2}=1"

"B=\frac{2}{85}"

"A=\frac{9}{85}"


Answer: particular solution is

"y_p=\frac{9}{85}e^xcosx+\frac{4}{85}e^xsinx"